LBT SW Testing and Range Balancing:
UT 2011 Jan 27

Observer: DThompson
Telescope Operator: DHuerta
Instrument Support: JMorris
Telescope Support: JUrban
SW support:

Summary

  • TCS BP8 testing
  • LUCI quick checkout
    • readnoise nominal
    • Mask FPU to TURNOUT took 93 seconds
    • N1.8 sieve mask images aberrated at bottom of detector
    • Took reference images in 1.0, 1.5, 2.0 arcsec slits
  • New LUCI pointing model with range-balanced collimation model (2.14" rms)

Details

First night back up after bogie fix. AZ max speed cut in half (0.65deg/sec).

00:15 Run up IRTC: wide-field, H filter 245K on the cooling.

00:40 LUCI engineer to restart TelescopeService since we switched to BP8. Problem getting WCS keywords, and it is still trying to sync 20110118. No images saved, no readout displayed on GEIRS display. Restarting GEIRSServer. JMorris reports lights in Data Interface box different than earlier today.

0056 Restart LUCI as observer. NXClient stops at "Negotiating link parameters" and times out. Stephen restarted NXServer and it still did not work. Odd, but it just started working for some reason (UT01:12). Still not able to read the detector and still having trouble reading WCS DD entries.

01:52 We shutdown MCE and ROE power and waited a few minutes, then stopped and restarted GeirsService, then did a clean_geirs and start_geirs and now we can read and save images. We still get the "Unable to get WCS keywords" error. Headers look OK except for temperature sensor 306 (reads 999999). The IMGUI came up blank (no green fields) - initialize gave some errors starting at 01:50:51 but the GUI comes up looking OK. Initialize again...seems OK and the WCS error is now gone.

InformationTechnology this for LUCI: The "Unable to get WCS keywords" error message appears to have been caused by the instrument not being initialized after a power cycle this afternoon. So, the log message was uninformative.

02:16 Glass is 9 degrees warner than outside! DGH is ventilating while the SW tests are done.

05:09 LUCI functional checkout
  • Check system time is same as our GPS ntp server (same)
  • Run /home/engineer/bin/showSupervisorStatus.sh (all looks OK)
  • Check LUCI1 Web-IO Control, verify that ROE and MCE are powered (they are on)
  • Start LUCIFER Control SW and check that all is running as needed (OK after initialize)
  • Open UI and check that telescope data are being updated (OK after restart)
  • Check darks (see: LUCIFER$/readnoise script)
    • O2DCR 2.0s (11.68 e-)
    • MER10 10.0s (5.44 e-)
  • Take sieve mask images in N3.75 (#0012) and N1.8 (#0013, 14) to check focus gradients (N3.75 looks OK, N1.8 has "comet tails" pointing up on almost everything in the lower half of the image)

05:30 Reconfigure for on-sky tests. IRTC pointing models on both sides plus the latest "range-balanced" collimation model on each side. Acquire source on IRTC and collimate, then check pointing with LUCI and adjust with mode 1 pointing changes until you can acquire the guide star on the LUCI side.

08:20 Continue with LUCI tests...
  • Move of mask from fpu to turnout too 93 seconds.
  • Move calibration unit in (unit moves in and out after doing it once manually on webcal)
    • Check that lamps work (lamps come on and this is noted in log, but the GUI indicator squares do not turn green)
    • Check gratings/tilts
    • Move calibration unit out (worked)

Image long-slits for reference
  • LS 600 0015
  • LS 900 0016
  • LS1200 0017

10:50 We are open! Tout=-4.7 (it was -8 a little while ago) Tmirror=-1.9

0019 ACT0277 at EL=76
0020 WT10_171 at EL=48

12:02 Pointing data, logging started in /home/telescope/PointingLogs/20110127_SX.log
LUCI# SourceID RA ...J2000... DEC
0027 WT10_408 20:38:05.102293818 +76:50:00.101170639
0028 WT10_404 08:45:56.803069314 +73:17:33.310262197
0029 ACT0337 11:13:54.561946813 +39:58:38.655706024
0030 WT10_271 08:24:45.406468395 +17:10:52.919420329
0031 WT10_241 12:13:29.526739891 +10:02:26.758246841
0032 ACT0354 14:41:05.993304934 +45:24:46.023968748
0033 ACT0186 11:12:46.411718445 -05:07:00.988237430
0034 ACT0088 12:49:32.497225607 -34:47:28.073272278
0035 WT10_175 15:47:38.045658147 -09:28:49.407059000
0036 ACT0232 17:59:27.078562695 +06:25:08.100508595
0037 WT10_330 17:13:27.222801237 +38:59:05.980864545
0038 WT10_294 13:23:46.711578900 +25:32:52.610855417
0039 WT10_300 17:52:32.177573728 +27:11:36.952862059
0040 WT10_317 18:38:58.561894795 +32:13:45.365475386
0041 ACT0264 15:44:48.123822624 +16:44:14.391811408
0042 ACT0209 14:35:59.031941077 +00:13:11.159368795

The resulting mode: LUCIFERSX.20110127.ptmod (2.14 arcsec rms):
LBT left Pointing Log, 2011 Jan 27 
T   13   2.1424    0.000   0.0000 
  IA       -359.4236     9.95554 
  IE       -115.7831     2.91548 
  NPAE      +39.1416    11.64814 
  AW         +1.9666     0.69639 
  AN        +25.5375     0.77444 
  HSCA2      -7.7039     1.12156
  HESE2     +93.0614     4.55603 
  CA        +20.4825    14.40470 
  TF        +84.9505     3.30752 
END 

Pointing model fit to 20110127 LUCI data:
tpoint.20110127.jpg

-- DavidThompson - 27 Jan 2011
I Attachment Action Size Date Who Comment
tpoint.20110127.jpgjpg tpoint.20110127.jpg manage 54 K 28 Jan 2011 - 01:31 DavidThompson Pointing model fit to 20110127 LUCI data
Topic revision: r8 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback